OpenStudy (baby456):

A teenager named Angie, wants a new pair of shoes. she saved up 30 dollars.she works at a store her hourly wage was 7.50,per hour. how long would she have to work if the shoes 123$. This is a Inequality probelm so please help.

OpenStudy (anonymous):

30 + 7.5x >= 123

OpenStudy (anonymous):

7.5x >= 123 - 30 x >= 93/7.5 x >= 12.4 13 hours

OpenStudy (baby456):

is that a greater than or equal too sign.

OpenStudy (anonymous):

yes